14-51 The solenoid used in an MRI scanner is 2 m long and 90 cm in diameter. It is wound froma copper wire 1.4 mm in diameter, with adjacent turns separated by an insulating layer of varnish with negligible thickness. See example 26.10 of your textbook. 4. Find the current that will produce a 0.5 Find the voltage that needs to be applied to generate this current. Resistivity of the copper is 1.75 x 10 m. Hint: find the length of the wire first 5. T magnetic field. a. 203 A b. 557 A с. 897 A d. 1234 A e. 2315A a. 15 kV b. 25 kV c. 35 kV d. 45 kV e. 55 kV
Answer #1

The length of the solenoid is, L = 2 m. And the diameter of the copper wire is, d = 1.4 mm. So, the no. of turns in the copper wire is

N = \frac {L} {d} =\frac{2}{1.4\times 10^{-3}}=1428

So, from the magnetic field formula in the solenoid, we write

7 1428 π×10

\Rightarrow I=\frac {2 \times 0.5} {4\pi \times 10^{-7}\times 1428}=557~A

So, option (b) is correct.

5)

Length of the wire is

l=N\times \pi D

Where D is the diameter of a single turn of the solenoid.

So, putting the given values, we get

1-1428 × π × 0.9 4037.57 772

The cross sectional area of the wire is

A = π /4 = π × ( 1.4 × 10-3)2 /4 = 1.539 × 10-0 m 7n

The resistance of the wire is

4037.57 1.539 × 10-6 = ρ = 1.75 × 10-8 ×

Voltage difference

V=IR =557\times 45.9=25.566~kV

So, option ( b) is correct.